OpenStudy (anonymous):

A 10.0 Ω resistor is in parallel with a 20.0 Ω resistor and supplied by a 9.00 V battery. What is the total current passing through the circuit?

OpenStudy (michele_laino):

it is simple the toatl resistance is: R=20+10=30 Ohm so the requested current is: I= V/R= 9/30=...Amperes

OpenStudy (michele_laino):

total*

OpenStudy (anonymous):

The answer choices are 0.38 A 0.75 A 1.35 A 6.00 A So do I go with 0.38 A

OpenStudy (michele_laino):

sorry I have made an error

OpenStudy (michele_laino):

the total resistance is: \[\Large R = \frac{1}{{\frac{1}{{10}} + \frac{1}{{20}}}} = ...Ohms\]

OpenStudy (michele_laino):

what is R?

OpenStudy (michele_laino):

hint: R=6.67 Ohms so the requested current, is: I=V/R= 9/6.67=...Amperes

OpenStudy (anonymous):

I got that answer but it doesn't match any answer choice.

OpenStudy (michele_laino):

what is your answer?

OpenStudy (anonymous):

oh I messed up I forgot to divide 9/6.67 I am so dumb the answer is 1.35A

OpenStudy (michele_laino):

the answer is 1.35 Amperes, good job!

OpenStudy (anonymous):

Thank you for your help. I should have been able to do this on my own. I just missed a tiny detail when I was trying to solve it myself.

OpenStudy (michele_laino):

thanks! :)